OpenStudy (anonymous):

6y=192 A. -32 B.-18 C.18 D.32

Vocaloid (vocaloid):

divide both sides by 6

OpenStudy (escamer):

A.

OpenStudy (escamer):

32

OpenStudy (alexandervonhumboldt2):

no direct answers

OpenStudy (escamer):

sorry, was just too tempting

OpenStudy (alexandervonhumboldt2):

first divide both sides by 6: 6y=192 6y/6=192/6 y=32

OpenStudy (alexandervonhumboldt2):

do what vocloid said

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so its D ?

OpenStudy (escamer):

oh yeah its D i didnt see the negative lol my bad

OpenStudy (anonymous):

ok thanks